If we have the following functions:
f(x) = 3x – 2 and g(x) = x2 – 2x find the following:
a. (f ○ g)(1)
b. (g ○ f)(1)
c. (f ○ g)(x)

Answer by eperette(173)

a) f(g(1))
g(1)=(1)^2 -2(1)
g(1)= 1 -2
g(1)=-1
f(g(1))=f(-1)
f(-1)=3(-1)-2
f(-1)=-3-2
f(-1)=-5
therefore f(g(1))=-5
***********************************
b) g(f(1))
f(1)=3(1)-2
f(1)=3-2
f(1)=1
g(f(1))=g(1)
g(1)=(1)^2-2(1)
g(1)=1-2
g(1)=-1
therefore g(f(1))=-1
************************************
f(g(x))=f(x^2-2x)
f(x^2-2x)= 3(x^2-2x)-2= 3x^2 - 6x -2